graph each exponential function
A.)f(x)=2*(1/2)
B.)k(x)=3*()1/3
C.)h(x)=3*(3)
D.)g(x)=5*(1.2)

Answer by Edwin McCravy(20055)

graph each exponential function
A.)f(x)=2*(1/2)
B.)k(x)=3*()1/3
C.)h(x)=3*(3)
D.)g(x)=5*(1.2)

You copied those wrong. There
must be an x on the right side
of those equations. Repost the
correct equations.
Edwin